the 888 mile trip to the mountains took three times as long as the 280 mile trip to the ocean. if the speed to the ocean was 4 mph less than the speed to the mountains, find both speeds and both times
Answer by mananth(16949)

Mountain distance =888 miles-- speed = x
time (mountain)= 888/x

ocean distance = 280 miles - let speed be x-4
time (ocean)= 280/(x-4)

Time ( mountain)= 3*time(ocean)
888/x= 3(280/(x-4)
cross multiply
888(x-4)=3*280*x
888x-3552=840x
48x=3552
x=3552/48
x= 74 mph => mountain speed
Ocean speed => 70-4=70 mph
m.ananth@hotmail.ca